OpenStudy (anonymous):

A rectangle is 6 times as long as it is wide. The perimeter is 90 cm. Find the dimensions of the rectangle

OpenStudy (anonymous):

P = 2L + 2W , where P is perimeter , L is length, W is width we also know that L = 6W P = 2(6W) + 2W P = 12W + 2W P = 14W , plug in values 90 = 14W W = 90/14 I'll let you find the length.

OpenStudy (anonymous):

cool thank you

OpenStudy (anonymous):

no prob